A body of mass 700g falls from a height of 10m.Find the momentum of the

body.


1
Expert's answer
2021-10-10T16:07:56-0400

"s=v^2/(2g)\to v=\sqrt{2gs}=\sqrt{2\cdot 9.8\cdot 10}=14\ (m/s)"


"p=mv=0.7\cdot14=9.8\ (kg\cdot m/s)" . Answer